find the focus of the parabola that has the vertex (-3,2), opens horizontally, and passes through the point, (-10,1).
I can get to 4p= -1/7 but get stuck there trying to get the focus.

Answer by rothauserc(4718)

Use the following formula for a horizontal parabola
4p(x-h) = (y-k)^2
the vertex is (-3,2) and parabola passes thru point (-10,1), therefore
4p(-10+3) = (1-2)^2
-28p = 1
p = -1/28 = -0.035714286 approx -0.04
p is the distance from the vertex to the focus, therefore
focus = (-3.04, 2)